Write an equation in slope intercept form of the line that passes through the given point and is parallel to the graph of the given equation.
(1,3) y= 3x+2

Answer by Alan3354(69443)

Write an equation in slope intercept form of the line that passes through the given point and is parallel to the graph of the given equation.
(1,3) y= 3x+2
================================================
Do it like this one:
The equation of a line parallel to the given line but passing through the given point
Y= 1/4x-2; (8,-1)
-------------
Slope m = 1/4
---
--> y+1 = (1/4)*(x-8)